# On the number of modes of Gaussian kernel density estimators

### v1

@@ 1.1 Setup and main result
For $\beta>0$ and $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$, the Gaussian kernel density estimator (KDE) with bandwidth $h=\beta^{-1/2}$ is defined as
$\widehat{P}_{n}(t):=\frac{1}{n}\sum_{i=1}^{n}\mathsf{K}_{h}\ast\delta_{X_{i}}(t)=\frac{\sqrt{\beta}}{n\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{\beta}{2}(t-X_{i})^{2}},\hskip 28.45274ptt\in\mathbb{R}.$ (1.1)
@@
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Asymptotically as $n\to\infty$, the expected number of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is $1+o(1)$ if $1\ll\beta\ll n^{2/3}$, and $\widetilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\beta}}\right)$ if $n^{2/3}\lesssim\beta\ll n^{2}/\log^{6}n$.
@@
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $\beta^{-1/2}$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Suppose $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, and 4.2, then asymptotically as $n,\beta\to\infty$,
@@
4.2 is related to the decay of the tails of the modulus of continuity of $\widehat{P}_{n}^{\prime\prime}(\cdot)$, and is needed to apply the Kac-Rice formula (Theorem 2.1), which states that the expected number of modes of $\widehat{P}_{n}$ is some conditional expectation with respect to the joint law of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$. We postpone a further discussion to Section 4.1.
@@
  • To better appreciate the range of values for $\beta$ in this theorem as well as subsequent ones, we use minimax theory as a benchmark; see, e.g., [Tsy09]. The reparametrization $h=\sqrt{\beta}$ is motivated by the connection to the Transformer model described in Section 1.2. Using an optimal bias-variance tradeoff [Tsy09, Chapter 1], we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$. This gives $\beta\asymp n^{\frac{2}{2s+1}}$. For $s\in(0,\infty)$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$. In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem 1.1 is achieved for $\beta\approx n^{2/3}$, which is the optimal choice for Lipschitz densities. The message of our main Theorem 1.2 below is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
@@
$T:=\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-\log\beta-\omega(\beta)},\sqrt{2\log n-\log\beta-\omega(\beta)}}\right]$ (1.2)
@@
$T^{\prime}:=\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}}\right]$ (1.3)
if $\beta\leqslant n^{2/3}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{2/3}$.
We see how Theorem 1.2 implies Theorem 1.1 with $h=\beta^{-1/2}$. From the former, we see that almost all the modes lie in $T\setminus T^{\prime}$. If $h\gg n^{-1/3}$ so that $\sqrt{2\log n-3\log\beta}\gg 1$, then $T\setminus T^{\prime}$ is disjoint from $[a,b]$, so there are few modes in $[a,b]$; if $h\ll n^{-1/3}$ so $\sqrt{2\log n-3\log\beta}\ll 1$, the fixed interval $[a,b]$ contains a near constant fraction of length of $T\setminus T^{\prime}$ and thus a near constant fraction of the modes.
@@ 1.2 Motivation
evolving on the unit sphere $\mathbb{S}^{d-1}$ because of $\mathsf{P}_{x}^{\perp}:=I_{d}-xx^{\top}$. Here, $\tau\geqslant 0$ plays the role of layers, the $n$ particles $x_{i}(\tau)$ represent tokens evolving through a dynamical system. This system is characterized by a temperature parameter $\beta\geqslant 0$ that governs the space localization of particle interations. One sees that all particles move in time by following the field $\nabla\log(\mathsf{K}_{\beta^{-1/2}}\ast\mu_{\tau})$; here, $\mu_{\tau}$ is the empirical measure of the particles $x_{1}(\tau),\ldots,x_{n}(\tau)$ at time $\tau$.
@@ 1.3 Sketch of the proof
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then under 4.2,
@@
To show the validity of the Gaussian approximation, we use the Edgeworth expansion of the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ around the Gaussian distribution with matching first two moments. We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [BR10, Theorem 19.2]. We note that [KM97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
@@ 2.1 The Kac-Rice formula
    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\to 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity111defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leq\eta}|f(t)-f(s)|$. of $\Psi^{\prime}(\cdot)$.
@@
$U_{u}(\Psi,T):=\mathopen{}\mathclose{{}\left|\{t\in T:\Psi(t)=u,\Psi^{\prime}(t)>0\}}\right|.$
@@ 2.2 Computing the Gaussian approximation
$\mu_{t}$ $:=\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ -t^{2}\end{bmatrix}$ (2.5)
$\Sigma_{t}$ $:=\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\asymp\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}1&-t\\ -t&\beta\end{bmatrix}.$
@@
$\varphi(x):=\frac{1}{\sqrt{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
@@
$\Omega:=\Sigma_{t}^{-1}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}\beta&t\\ t&1\end{bmatrix}.$
@@ 2.3 The Kac-Rice integral over $\varphi$
As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by 1.2. Let $t_{s}:=\sqrt{2\log n-s\log\beta}$. As the integrand is positive, for constants $C,C^{\prime}>0$
@@ 3.1 Bounding the third order error
$\psi(\mathbf{x}):=\frac{\varphi(\mathbf{x})}{6}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}(\mathbf{x})\quad\text{where}\quad H^{\alpha}:=(-1)^{|\alpha|}\frac{\partial^{\alpha}\varphi}{\varphi}$ (3.2)
@@
$\eta_{3}:=\max\mathopen{}\mathclose{{}\left\{\kappa_{t}^{(k,3-k)}:k\in\{0,1,2,3\}}\right\}\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}.$
We bound $H^{(k,3-k)}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-1/2}[(0,y)-\mu_{t}]}\right)$ by a polynomial in $\tilde{y}=\Theta(\beta^{1/2}e^{{t^{2}}/{4}}y)$. Now, by a similar method as Lemma 2.3, we obtain the following bound. It says that when we integrate the Edgeworth series $p_{t}=\varphi+n^{-1/2}\psi+\dots$ over $y$ and $t$, the contribution $\varphi$ dominates $n^{-1/2}\psi$, hinting at the validity of the approximation.
@@ 3.2 Bounding higher order errors
We defer the discussion of the assumption to 4.2 and defer the proof to Section A.3. Here, we comment on the differences with [BR10, Theorem 19.2] for $s=3$, which we follow in spirit but modify to allow $n$ dependence in $q_{t}$ in the form of $\beta$. There, it is shown that the left hand side is $o(n^{-1/2})$ provided third moments $\eta_{3}=O(1)$. With the bound 3.1 on $\eta_{3}$ in our case which is not constant, the error from the asymptotic series decays not in powers of $n^{-1/2}$ but powers of $O(n^{-1/2}e^{{t^{2}}/{4}}\beta^{1/4})=O(\exp(-\omega(\beta)/2))$.
@@
Let $\beta,n$ be as in Theorem 1.2, and fix $t\in T$. Let $\upmu_{t}$ denote the law of $(F_{n}(t),F_{n}^{\prime}(t))$ defined in (2.2). Then $\upmu_{t}$ admits a density $p_{t}\in\mathscr{C}^{0}(\mathbb{R}^{2})$ satisfying $p_{t}({\bf{x}})\to 0$ as $\|\bf{x}\|\to\infty$. Moreover, conditions 1, 2 in Theorem 2.1 also hold for $\Psi(t)=F_{n}(t)$.
We defer the proof to Section A.4. We work under assumption
Assumption 4.2.
Consider the setting of Proposition 4.1. We assume that condition 4 of Theorem 2.1 holds for $\Psi(t)=F_{n}(t)$.
To check conditions on moduli of continuity such as 4 in Theorem 2.1 in the Gaussian setting, one usually resorts to using results such as the Borell-TIS inequality [AT09, Theorem 2.1.1]. Checking the validity of this assumption in the present, non-Gaussian, setting does not appear straightforward.
With Proposition 4.1 and 4.2, we deduce
@@
$g_{i}(t):=\sqrt{\frac{\beta}{2\pi n^{2}}}\mathsf{K}_{\beta^{-1/2}}\mathopen{}\mathclose{{}\left(t-X_{i}}\right).$ (4.2)
@@ Appendix A Additional proofs
  • $\mathbb{E}G^{2}\asymp\beta^{-\frac{5}{2}}(t^{2}+\beta)e^{\frac{t^{2}}{2}}\gg\beta^{-3}t^{2}E^{2}_{t}\asymp(\mathbb{E}G)^{2},$
  • $\mathbb{E}GG^{\prime}\asymp-\beta^{-\frac{7}{2}}e^{\frac{t^{2}}{2}}(\beta^{2}t+\beta t^{3})\gg\beta^{-4}E^{2}_{t}t(\beta-\beta t^{2})\asymp(\mathbb{E}G)(\mathbb{E}G^{\prime}),$
  • $\mathbb{E}G^{\prime 2}\asymp\beta^{-\frac{9}{2}}e^{\frac{t^{2}}{2}}(\beta^{4}+\beta^{2}t^{4})\gg\beta^{-5}E^{2}_{t}\beta^{2}(1+t^{4})\asymp(\mathbb{E}G^{\prime})^{2},$
where we bound $e^{\frac{t^{2}}{2}}\leqslant 1$. From these computations, and the remark after A.1, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in 2.2.∎
@@
$\delta:=\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}(\mathbf{z})}\right|<1.$
@@
Point 1 in Theorem 2.1 can readily be seen to hold because of the explicit form of both of the fields. We focus on showing Point 3, the proof of which can be repeated essentially verbatim to deduce Point 2.
@@
$=\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|^{n}\,\mathrm{d}\xi.$
@@
We can conclude by the Lebesgue dominated convergence theorem.∎

### v2

@@ 1.1 Setup and main result
For $\beta>0$ and $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$, the Gaussian kernel density estimator (KDE) with bandwidth $h=\beta^{-\frac{1}{2}}$ is defined as
$\widehat{P}_{n}(t)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathsf{K}_{h}\ast\delta_{X_{i}}(t)=\frac{\sqrt{\beta}}{n\sqrt{2\pi}}\sum_{i=1}^{n}e^{-\frac{\beta}{2}(t-X_{i})^{2}},\hskip 28.45274ptt\in\mathbb{R}.$ (1.1)
@@
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Asymptotically as $n\to\infty$, the expected number of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is $1+o(1)$ if $1\ll\beta\ll n^{\frac{2}{3}}$, and $\widetilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\beta}}\right)$ if $n^{\frac{2}{3}}\lesssim\beta\ll n^{2}/\log^{6}n$.
@@
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $\beta^{-\frac{1}{2}}$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Suppose $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$.Then asymptotically as $n,\beta\to\infty$,
@@
Several comments are in order.
@@
  • To better appreciate the range of values for $\beta$ in this theorem as well as subsequent ones, we use minimax theory as a benchmark; see, e.g., [Tsy09]. The reparametrization $h=\beta^{-\frac{1}{2}}$ is motivated by the connection to the Transformer model described in Section 1.2. Using an optimal bias-variance tradeoff [Tsy09, Chapter 1], we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$. This gives $\beta\asymp n^{\frac{2}{2s+1}}$. For $s>0$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$. In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem 1.1 is achieved for $\beta\approx n^{\frac{2}{3}}$, which is the optimal choice for Lipschitz densities. The message of our main Theorem 1.2 below is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
@@
$T\coloneqq\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-\log\beta-\omega(\beta)},\sqrt{2\log n-\log\beta-\omega(\beta)}}\right]$ (1.2)
@@
$T^{\prime}\coloneqq\mathopen{}\mathclose{{}\left[-\sqrt{2\log n-3\log\beta},\sqrt{2\log n-3\log\beta}}\right]$ (1.3)
if $\beta\leqslant n^{\frac{2}{3}}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{\frac{2}{3}}$.
We see how Theorem 1.2 implies Theorem 1.1 with $h=\beta^{-\frac{1}{2}}$. From the former, we see that almost all the modes lie in $T\setminus T^{\prime}$. If $h\gg n^{-\frac{1}{3}}$ so that $\sqrt{2\log n-3\log\beta}\gg 1$, then $T\setminus T^{\prime}$ is disjoint from $[a,b]$, so there are few modes in $[a,b]$; if $h\ll n^{-\frac{1}{3}}$ so $\sqrt{2\log n-3\log\beta}\ll 1$, the fixed interval $[a,b]$ contains a near constant fraction of length of $T\setminus T^{\prime}$ and thus a near constant fraction of the modes.
@@ 1.2 Motivation
evolving on the unit sphere $\mathbb{S}^{d-1}$ because of $\mathsf{P}_{x}^{\perp}\coloneqq I_{d}-xx^{\top}$. Here, $\tau\geqslant 0$ plays the role of layers, the $n$ particles $x_{i}(\tau)$ represent tokens evolving through a dynamical system. This system is characterized by a temperature parameter $\beta\geqslant 0$ that governs the space localization of particle interations. One sees that all particles move in time by following the field $\nabla\log(\mathsf{K}_{\beta^{-1/2}}\ast\mu_{\tau})$; here, $\mu_{\tau}$ is the empirical measure of the particles $x_{1}(\tau),\ldots,x_{n}(\tau)$ at time $\tau$.
@@ 1.3 Sketch of the proof
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then
@@
To show the validity of the Gaussian approximation, we use the Edgeworth expansion of the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ around the Gaussian distribution with matching first two moments. We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [BR10, Theorem 19.2]. This strategy has been used in [BCP19], in a different context. We note that [KM97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
@@ 2.1 The Kac-Rice formula
    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\searrow 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity111defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leqslant\eta}|f(t)-f(s)|$. of $\Psi^{\prime}(\cdot)$.
@@
$U_{u}(\Psi,T)\coloneqq\mathopen{}\mathclose{{}\left|\{t\in T:\Psi(t)=u,\Psi^{\prime}(t)>0\}}\right|.$
@@ 2.2 Computing the Gaussian approximation
$\mu_{t}$ $\coloneqq\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ -t^{2}\end{bmatrix}$ (2.5)
$\Sigma_{t}$ $\coloneqq\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\asymp\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}1&-t\\ -t&\beta\end{bmatrix}.$
@@
$\varphi(x)\coloneqq\frac{1}{\sqrt{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
@@
$\Omega\coloneqq\Sigma_{t}^{-1}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}\beta&t\\ t&1\end{bmatrix}.$
@@ 2.3 The Kac-Rice integral over $\varphi$
As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by 1.2. Let $t_{s}\coloneqq\sqrt{2\log n-s\log\beta}$. As the integrand is positive, for constants $C,C^{\prime}>0$
@@ 3.1 Bounding the third order error
$\psi(\mathbf{x})\coloneqq\frac{\varphi(\mathbf{x})}{6}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}(\mathbf{x})\quad\text{where}\quad H^{\alpha}\coloneqq(-1)^{|\alpha|}\frac{\partial^{\alpha}\varphi}{\varphi}$ (3.2)
@@
$\eta_{3}\coloneqq\max\mathopen{}\mathclose{{}\left\{\kappa_{t}^{(k,3-k)}:k\in\{0,1,2,3\}}\right\}\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}.$
We bound $H^{(k,3-k)}\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-1/2}[(0,y)-\mu_{t}]}\right)$ by a polynomial in $\tilde{y}=\Theta(\beta^{1/2}e^{{t^{2}}/{4}}y)$. Now, by a similar method as Lemma 2.3, we obtain the following bound. It says that when we integrate the Edgeworth series $p_{t}=\varphi+n^{-\frac{1}{2}}\psi+\dots$ over $y$ and $t$, the contribution $\varphi$ dominates $n^{-\frac{1}{2}}\psi$, hinting at the validity of the approximation.
@@ 3.2 Bounding higher order errors
We defer the proof to Section A.3. Here, we comment on the differences with [BR10, Theorem 19.2] for $s=3$, which we follow in spirit but modify to allow $n$ dependence in $q_{t}$ in the form of $\beta$. There, it is shown that the left hand side is $o(n^{-\frac{1}{2}})$ provided third moments $\eta_{3}=O(1)$. With the bound 3.1 on $\eta_{3}$ in our case which is not constant, the error from the asymptotic series decays not in powers of $n^{-\frac{1}{2}}$ but powers of $O(n^{-1/2}e^{{t^{2}}/{4}}\beta^{1/4})=O(\exp(-\omega(\beta)/2))$.
@@
Let $\beta,n$ be as in Theorem 1.2, and fix $t\in T$. Let $\upmu_{t}$ denote the law of $(F_{n}(t),F_{n}^{\prime}(t))$ defined in (2.2). Then $\upmu_{t}$ admits a density $p_{t}\in\mathscr{C}^{0}(\mathbb{R}^{2})$ satisfying $p_{t}({\bf{x}})\to 0$ as $\|\bf{x}\|\to\infty$. Moreover, conditions 1, 2, 4 in Theorem 2.1 also hold for $\Psi(t)=F_{n}(t)$, thus Theorem 2.1 applies.
We defer the proof to Section A.4. With Proposition 4.1, we deduce
@@
$g_{i}(t)\coloneqq\sqrt{\frac{\beta}{2\pi n^{2}}}\mathsf{K}_{\beta^{-1/2}}\mathopen{}\mathclose{{}\left(t-X_{i}}\right).$ (4.2)
@@ Appendix A Additional proofs
  • $\mathbb{E}G^{2}\asymp\beta^{-\frac{5}{2}}(t^{2}+\beta)e^{-\frac{t^{2}}{2}}\gg\beta^{-3}t^{2}e^{-t^{2}}\asymp(\mathbb{E}G)^{2},$
  • $\mathbb{E}GG^{\prime}\asymp-\beta^{-\frac{7}{2}}e^{-\frac{t^{2}}{2}}(\beta^{2}t+\beta t^{3})\gg\beta^{-4}e^{-t^{2}}t(\beta-\beta t^{2})\asymp(\mathbb{E}G)(\mathbb{E}G^{\prime}),$
  • $\mathbb{E}G^{\prime 2}\asymp\beta^{-\frac{9}{2}}e^{-\frac{t^{2}}{2}}(\beta^{4}+\beta^{2}t^{4})\gg\beta^{-5}e^{-t^{2}}\beta^{2}(1+t^{4})\asymp(\mathbb{E}G^{\prime})^{2},$
where we bound $e^{-\frac{t^{2}}{2}}\leqslant 1$. From these computations, and the remark after A.1, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in 2.2.∎
@@
$\delta\coloneqq\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}(\mathbf{z})}\right|<1.$
@@
Point 1 in Theorem 2.1 can readily be seen to hold because of the explicit form of both of the fields. Point 4 also readily holds, since $F_{n}^{\prime\prime}$ is a Lipschitz function for every realization of $X_{i}$, as a sum of Lipschitz functions. We focus on showing Point 3, the proof of which can be repeated essentially verbatim to deduce Point 2.
Proof of Point 3
@@
$=\frac{1}{(2\pi)^{\frac{n}{2}}}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|^{n}\,\mathrm{d}\xi.$
@@
We can conclude by the Lebesgue dominated convergence theorem.